# Project Page: https://yongtaoge.github.io/projects/humanwild

### v2

@@ Abstract
Despite remarkable progress has been made on the problem of 3D human pose and shape estimation (HPS), current state-of-the-art methods rely heavily on either confined indoor mocap datasets or datasets generated by rendering engine using computer graphics (CG). Both categories of datasets exhibit inadequacies in furnishing adequate human identities and authentic in-the-wild background scenes, which are crucial for accurately simulating real-world distributions. In this work, we show that synthetic data created by generative models is complementary to CG-rendered data for achieving remarkable generalization performance on diverse real-world scenes. Specifically, we propose an effective approach based on recent diffusion models, termed $\tt HumanWild$, which can effortlessly generate human images and corresponding 3D mesh annotations. We first collect a large-scale human-centric dataset with comprehensive annotations, e.g., text captions and surface normal images. Then, we train a customized ControlNet model upon this dataset to generate diverse human images and initial ground-truth labels. At the core of this step is that we can easily obtain numerous surface normal images from a 3D human parametric model, e.g., SMPL-X, by rendering the 3D mesh onto the image plane. As there exists inevitable noise in the initial labels, we then apply an off-the-shelf foundation segmentation model, i.e., SAM, to filter negative data samples. Our data generation pipeline is flexible and customizable to facilitate different real-world tasks, e.g., ego-centric scenes and perspective-distortion scenes. By exclusively employing generative models, we generate large-scale in-the-wild human images and high-quality annotations, eliminating the need for real-world data collection. The generated dataset comprises 0.79M images with corresponding 3D annotations, covering versatile viewpoints, scenes, and human identities. We train various HPS regressors on top of the generated data and evaluate them on a wide range of benchmarks (3DPW, RICH, EgoBody, AGORA, SSP-3D) to verify the effectiveness of the generated data.
Keywords:
Controllable Human Generation Synthetic Dataset 3D Human Reconstruction Diffusion Model
@@ 1 Introduction
Estimating human pose and shape (HPS) [27, 41, 36, 39, 11] from a single RGB image is a core challenge in computer vision and has many applications in robotics [10, 82, 75], computer graphics [90, 74], and digital content creation [76]. Current HPS estimation methods require well-annotated datasets to achieve good performance. Unfortunately, collecting large-scale 3D human body data is time-consuming and expensive.
@@
As shown in Tab. 1, contemporary methodologies for acquiring precise 3D human body data predominantly employ two primary pipelines. The initial pipeline encompasses indoor motion capture (mocap) systems, such as marker-based and vision-based systems, utilized by numerous existing datasets [22, 71, 8, 49] for human body attribute capture. Nonetheless, this mocap pipeline confronts two primary limitations: firstly, the intricate and costly nature of mocap systems, involving synchronization and operation complexities necessitating specialized expertise; secondly, the restricted number of actors in datasets, typically set against indoor or laboratory backgrounds, precluding large-scale human data collection in diverse settings. The alternative pipeline involves synthesizing 3D human datasets via CG rendering [5, 84, 79]. However, this approach also presents twofold challenges: firstly, the collection costs incurred in acquiring high-quality 3D assets, including avatars and scene elements, alongside the requisite expertise in 3D rendering; secondly, ensuring realism in synthetic human bodies and background scenes, despite persistent appearance domain gaps between rendered and real-world images. This domain gap is discernible in Fig. 1. where we visualize feature distributions extracted by DINO-v2 [53] through UMAP dimension reduction [47], with red points denoting real-world in-the-wild data and blue points representing CG-rendered data. Evidently, a conspicuous domain gap exists between real-world and CG-rendered data.
Since both mocap-based and CG-rendered datasets fail to provide diverse, complex in-the-wild images, some researchers have considered capturing pseudo 3D ground truth by estimating from 2D clues or using additional sensors. SMPLify [6] proposed to fit the parameters of a 3D human model to the location of 2D keypoints. EFT [25] introduced the Exemplar Fine-Tuning strategy by overfitting a pre-trained 3D pose regressor with 2D keypoint reprojection loss, taking the final output of the regressor as pseudo labels. Nevertheless, these methodologies encounter challenges in accurately generating camera parameters and body parameters, resulting in suboptimal performance on 3D human pose estimation benchmarks.
Contemporary generative models, e.g., Stable Diffusion XL [58] are trained on billion-scale text-image pairs. This extensive training regime has demonstrated efficacy in faithfully simulating real-world distributions. In this paper, we focus on creating a new automatic and scalable synthetic data generation pipeline solely through generative models. The challenge of the pipeline lies in the sampling of diverse data pairs from the prior distribution inherent to generative models. It’s crucial to ensure diversity in pose, shape, and scene representation within generated human images, which is paramount for accurately simulating real-world human distribution. Conversely, ensuring precise alignment between human images and generated annotations stands as another critical facet. This alignment is indispensable for facilitating the effective training of downstream tasks. DiffusionHPC [78] proposes a simple solution by providing diverse text prompts to the model and subsequently utilizing pre-trained 3D human pose estimation (HPS) models to obtain pseudo labels. Nonetheless, relying solely on textual prompts lacks the necessary granularity to precisely control aspects such as pose, shape, and spatial positioning of human bodies. Furthermore, the resultant 3D pseudo labels often exhibit significant levels of noise. In contrast, we propose a method that begins by sampling SMPL-X parameters, which represent human body configurations, from large-scale human motion capture datasets such as AMASS [44, 5]. Subsequently, we utilize a random camera setup to render the human mesh into a normal map, thereby adding an extra input condition. This normal map provides supplementary information about the surface orientation of the human body. Finally, we feed text prompts and normal maps to a customized ControlNet [88] for generating human images. Notably, we collect over 1M image-text-normal pairs from LAION5B [64] and MOYO-700M [71], Hi4D [86], and BEDLAM [5] to train the normal-based ControlNet. As such, our method attains precise control over the visual characteristics of generated human images while simultaneously acquiring initial training data pairs derived from the input conditions and resultant human images. Empirical investigations have uncovered instances of label noise within initial training data pairs, including occurrences where the generated human and input conditions form mirror pairs, or where the human head orientation in the image diverges from the input SMPL-X parameters. To mitigate this challenge, we leverage a pre-trained segmentation foundation model, namely SAM [30], to forecast the human mask of the generated images. Subsequently, we compute the intersection-over-union (mIoU) between the ground-truth human mask and the predicted human mask, thereby enabling the filtration of data samples with an IoU below a predetermined threshold.
@@
With the aforementioned pipeline, we can finally generate a large-scale 3D human dataset in the wild, with around 0.79M samples at $768\times 768$ resolution. Compared to previous datasets, our pipeline can generate diverse human identities and various in-the-wild scenes. Notably, the pipeline is much cheaper than both mocap-based and CG-based counterparts and is scalable to generate 3D human datasets in the wild with versatile human identities and real-world scenes.
Table 1: Comparison of different types of 3D HPS datasets.
Data type Datasets Subjects Scene Frame Modalities RGB D/N K2D K3D B.P. WB.P. Real-world Monocular COCO [42] - - 104K MPI-INF-3DHP [49] 8 - 1.4M MPII [1] - - 24K Multi-view HuMMan 1000 1 60M ZJU Mocap [57] 6 1 >1K AIST++ [38] 30 1 10.1M MOCAP 3DPW [46] 5 <60 51K Human3.6M [22] 11 1 3.6M Synthetic AGORA [55] >350 - 18K Synbody [84] 10K 6 1.2M BEDLAM [5] 217 103 280K Generated HumanWild $\infty$ $\infty$ 0.79M
@@
Our contributions can be summarised as follows. 1). We propose a fully automatic pipeline to synthesize realistic and diverse human images with well-aligned annotations, including SMPL-X parameters and text descriptions. The dataset can empower a wide range of downstream perception tasks by rendering SMPL-X mesh into corresponding annotation format, e.g., human pose and shape estimation, human part segmentation, and human surface normal estimation. 2). We verify the quality of the generated dataset on the 3D HPS task. Experimental results indicate that the proposed pipeline is compositional with CG-rendered data, enhancing performance across multiple challenging HPS benchmarks under consistent settings.
2 Related Works
@@ 2.1 Human Pose and Shape Estimation Datasets
Synthetic human pose datasets, developed with computer graphic techniques, has been used for many years. SURREAL [72] applies human skin and cloth textures to bare SMPL meshes, which lack realistic details. AGORA [55] uses high-quality static human scans for image rendering, but this routine also suffers from a high workload of scanning and rigging. However, rendering realistic manipulatable synthetic human datasets involves many challenges, including the need for diverse virtual properties for realistic data. BEDLAM [5] and Synbody [84] augment SMPL-X meshes [56] with diverse hair models and skin textures, facilitating the simulation of physically realistic cloth and hair dynamics. These processes can be resource-intensive. Furthermore, the use of rendering engines demands many professional skills. Thus, the rendering process can be computationally expensive and time-consuming.
Controllable human image generation has gained greate traction with the advancement of Stable Diffusion [60, 88]. Text2Human [24] uses a diffusion-based transformer sampler in response to text prompts and predicts indices from a hierarchical texture-aware codebook to conditionally generate realistic human images. HumanSD [26] introduces a skeleton-guided diffusion model with a novel heatmap loss for pose-conditioned human image generation. HyperHuman [43] proposes to jointly denoise surface normal and depth along with the synthesized RGB image conditioned on text prompt and pose skeleton.
Generative models for perception tasks is a promising field, which aims to enhance the performance of perception models with the capability of generative models. Several pioneer works have proposed to generate perception datasets with diffusion models. For instance, [2] and [73] verified the effectiveness of datasets synthesized by generative models in fundamental perception tasks, i.e., image classification, and object detection. StableRep [69] argues that training modern self-supervised methods on synthetic images from Stable Diffusion Models can yield impressive results. The learned representations often surpass those learned from real images of the same sample size. DatasetDM [80] leverages the prior of stable diffusion model by training customized perception decoders upon the output of the UNet [61], and then employs it to generate pseudo labels for semantic segmentation, depth estimation, and 2D human pose estimation.
@@ 3 Method
Figure 2: The overall pipeline of the proposed controllable data generation. Our ControlNet could be conditioned on fine-grained normal maps and structual text prompts. Our text prompt includes human appearance, pose, and indoor/outdoor scene types based on w. or w/o background normals.
We present $\tt HumanWild$, a simple yet effective approach for creating versatile human body images and corresponding perception annotations in a fully automated fashion, which can be used for many downstream human perception tasks, such as 2D/3D human pose and shape estimation (see Fig. 2). The core idea of the proposed pipeline is creating large-scale image-mesh-caption pairs by incorporating 2D generative models, e.g., ControlNet [88] and 3D human parametric models [56]. For the sake of completeness, we give a brief review of the controllable text-to-image (T2I), image-to-image (I2I) generative models and the 3D human parametric model, SMPL-X [56] in Sec. 3.1. In the following subsections, we illustrate how we generate the initial human image-annotation pairs in Sec. 3.2 and Sec. 3.3. Finally, we show how we filter out the noisy labels to get high-quality training pairs in Sec. 3.4.
@@ 3.1 Prerequisites
Stable Diffusion Models [60, 58] are text-to-image diffusion models capable of generating near photo-realistic images given any text input. It consists of an autoencoder and a U-Net. During training, U-Net model is designed to take a textual description and noise map $\epsilon$ as input, learning a forward noising process by gradually transforming an image latent variable, denoted as $\bar{y}\in\mathbb{R}^{h\times w\times c}$ into a noise map $\bar{y_{t}}\in\mathbb{R}^{h\times w\times c}$:
@@
where $t=[1,\cdots,T]$ is the timestep for controlling the noise level and ${\alpha}_{t}$ is the noise scheduler [19], $h$, $w$ and $c$ are the height, width, and channels of the latent variable. During inference, UNet takes a noise map as input and generates an approximation of the original latent variable using a denoise scheduler [67].
@@
SMPL-X [56], defined as $M(\boldsymbol{\beta},\boldsymbol{\theta},\boldsymbol{\psi}):\mathbb{R}^{|\theta|\times|\beta|\times|\psi|}\rightarrow\mathbb{R}^{3N}$, is a 3D wholebody human parametric model, employing shape $\boldsymbol{\beta}\in\mathbb{R}^{200}$, expression $\boldsymbol{\psi}\in\mathbb{R}^{50}$, and pose $\boldsymbol{\theta}\in\mathbb{R}^{55\times 3}$ to control the entire body mesh. The function of SMPL-X provides a differentiable skinning process that uses pose, shape, and expression parameters as inputs and delivers a triangulated mesh $V\in\mathbb{R}^{N\times 3}$ with $N=10475$ vertices. The reconstructed 3D joints $J\in\mathbb{R}^{144\times 3}$ can be obtained using a forward kinematics process. We use the SMPL-X model to parameterize the human body.
3.2 Surface Normal Estimation
The training of the ControlNet requires large-scale in-the-wild image and surface normal map pairs. However, it is impossible to manually annotate such data. Previous work [43] employed an off-the-shelf method, Omnidata [12] for generating pseudo surface normal maps. Unfortunately, in our pilot studies, we observed that surface normal maps generated by Omnidata [12] tend to lose details. Inspired by recent works that leverage diffusion prior for downstream perception tasks [28, 34, 80], we fine-tune a pretrained latent diffusion model for surface normal estimation. This refinement results in a deterministic mapping between input images and output surface normal maps. Drawing from strategies proposed in [3, 17], we introduce a blending strategy and redesign the diffusion process. Specifically, we treat the input image $x$ as the noise map and surface normal map $n$ as the generation target. We then extract corresponding latent variables $\bar{x}$ and $\bar{n}$ through the pre-trained VAE encoder. Finally, we reformulate the diffusion process described in Eq. 1 as:
$\bar{n_{t}}=\sqrt{\bar{\alpha}_{t}}\bar{n}+\sqrt{1-\bar{\alpha}_{t}}\bar{x}\quad t=[1,\cdots,T].$ (2)
See  Fig. 3 for illustration. To predict surface normals from a pre-trained T2I model, we fine-tune a UNet using the v-prediction objective [62] on a synthetic dataset, Hypersim [59]. During inference, the UNet progressively denoises input images and extracts surface normals using the DDIM scheduler [67].
Figure 3: The overall training pipeline of the surface normal estimation.

### v3

# Diffusion Models are Efficient Data Generators for Human Mesh Recovery

@@ Abstract
Despite remarkable progress having been made on the problem of 3D human pose and shape estimation (HPS), current state-of-the-art methods rely heavily on either confined indoor mocap datasets or datasets generated by a rendering engine using computer graphics (CG). Both categories of datasets exhibit inadequacies in furnishing adequate human identities and authentic in-the-wild background scenes, which are crucial for accurately simulating real-world distributions. In this work, we show that synthetic data created by generative models is complementary to CG-rendered data for achieving remarkable generalization performance on diverse real-world scenes. We propose an effective data generation pipeline based on recent diffusion models, termed HumanWild, which can effortlessly generate human images and corresponding 3D mesh annotations. Specifically, we first collect a large-scale human-centric dataset with comprehensive annotations, e.g., text captions, the depth map, and surface normal images. To generate a wide variety of human images with initial labels, we train a customized, multi-condition ControlNet model. The key to this process is using a 3D parametric model, e.g., SMPL-X, to easily create precise 2D keypoints, depth maps, and surface normal images by rendering the 3D mesh with specific camera parameters. As there exists inevitable noise in the initial labels, we apply an off-the-shelf foundation segmentation model to filter negative data samples, and a 2D vertex estimator to rectify the SMPL-X parameters by using SMPLify. Our data generation pipeline is both flexible and customizable, making it adaptable to various real-world tasks, such as human interaction in complex scenes and humans captured by wide-angle lenses. By relying solely on generative models, we can produce large-scale, in-the-wild human images with high-quality annotations, significantly reducing the need for manual image collection and annotation. The generated dataset encompasses a wide range of viewpoints, environments, and human identities, ensuring its versatility across different scenarios. To verify the effectiveness of the generated data, we perform comprehensive data ablation experiments by performing data ablation studies on top of the both generated data and existing datasets, and evaluating on a wide range of HPS benchmarks. We hope that our work could pave the way for scaling up 3D human recovery to in-the-wild scenes.
Index Terms:
Controllable Human Generation, and Synthetic Dataset, and Human Pose and Shape Estimation, and Diffusion Models
@@ 1 Introduction
Estimating human pose and shape (HPS) [1, 2, 3, 4, 5] from a single RGB image is a core challenge in computer vision and has many applications in robotics [6, 7, 8], computer graphics [9, 10], and digital content creation [11]. Current HPS estimation methods require well-annotated datasets to achieve good performance. Unfortunately, collecting large-scale, versatile 3D human body data is time-consuming and expensive.
Contemporary methodologies for acquiring precise 3D human body data predominantly employ two primary pipelines. The first pipeline involves indoor motion capture (mocap) systems, including both marker-based and vision-based approaches, which are employed by many existing datasets [12, 13, 14, 15] to capture human body attributes. However, this mocap pipeline faces two primary limitations. First, mocap systems are intricate and costly, and their complex operation demands specialized expertise. Second, the data captured is not very diverse. Datasets usually feature a limited number of actors in a lab or studio, which restricts the collection of large-scale human data in more varied environments. To diversify the data sources, some researchers also capture 3D pseudo ground truth by estimating from 2D clues or using additional sensors. SMPLify [16] proposed to fit the parameters of a 3D human model to the location of 2D keypoints. EFT [17] introduced the Exemplar Fine-Tuning strategy by overfitting a pre-trained 3D pose regressor with 2D keypoint reprojection loss, taking the final output of the regressor as pseudo labels. However, these methodologies face challenges in accurately generating camera parameters and body parameters, leading to sub-optimal performance on in-the-wild 3D human pose and shape estimation.
@@
An alternative approach is to generate 3D human datasets using computer-generated (CG) rendering, as seen in works like GTA-Human [18], BEDLAM [19] and SynBody [20]. However, this method has two main challenges. First, there are significant costs involved in obtaining high-quality 3D assets, such as avatars and scene elements, as well as the need for specialized skills in 3D rendering. Second, making synthetic humans and backgrounds look realistic is a major hurdle. There’s often a noticeable visual difference between the rendered images and real-world photographs. While Black et al. [19] show the potential of high-quality synthetic data to replace real images, they also highlight important limitations. Their work demonstrates that CLIFF [4], when trained only on synthetic data, can outperform models trained on real datasets. Conversely, they report that training HPS models from scratch with synthetic data, without leveraging pre-training on real-world 2D keypoints datasets (like COCO [21]), leads to inferior results. The study further reveals the significant positive impact of realistic clothing simulation compared to more basic textured bodies. This suggests that while synthetic data is a powerful tool, the degree of realism in the data is crucial for advancing 3D human pose and shape estimation.
While 2D keypoint detectors generalize well when trained on large in-the-wild datasets like COCO, significant gaps exist with 3D datasets. To investigate these gaps, we visualized the feature distributions of several key datasets. We extracted features from cropped human images using DINO-v2 and used UMAP for dimensionality reduction, as shown in Fig.˜1. Our analysis compares the real-world 2D datasets COCO [21] and MPII [22] against three distinct 3D datasets: the real-world outdoor 3DPW [23], the indoor Mocap MOYO [13], and the synthetically rendered BEDLAM [19]. The visualization reveals three key insights: (1). The feature distributions of the two real-world 2D datasets, COCO and MPII, are nearly identical. (2). In contrast, both the outdoor 3DPW and indoor MOYO datasets show distributions that diverge significantly from COCO. (3). The CG-rendered BEDLAM dataset is closer to COCO’s distribution than the other 3D datasets, though a noticeable gap remains.
This work introduces an automatic and scalable pipeline for generating synthetic data for 3D human pose and shape estimation. Our method uses generative models to create data with more realistic human distributions, overcoming a key limitation of current CG-rendered datasets, that diverse, high-quality CG assets do not scale easily. The intuition is that modern large generative models, e.g., Stable Diffusion XL [24] and FLUX [25] are trained on billion-scale text-image pairs. The extensive training regime has shown its efficacy in faithfully simulating real-world distributions. The challenge of the pipeline lies in the sampling of diverse data pairs from the prior distribution inherent to generative models. As analyzed in GTA-Human [18], ensuring diversity in camera angles, poses, body shapes, and human appearances within generated human images is critical to accurately simulate real-world human distributions. This diversity is essential for improving the performance of human pose and shape (HPS) tasks by providing more representative and varied training data, especially in-the-wild data. In contrast, ensuring precise alignment between human images and generated annotations is another critical aspect. This alignment is indispensable to facilitate the effective training of downstream tasks. DiffusionHPC [26] offers a straightforward solution by providing the model with diverse text prompts and using pretrained 3D human pose estimation (HPS) models to generate pseudo-labels. However, relying solely on text prompts lacks the granularity needed to precisely control important factors like pose, body shape, and spatial positioning. As a result, the generated 3D pseudo-labels often contain substantial noise, reducing their accuracy.
In contrast, we propose a generative method that leverages precise 3D annotations as conditions. We begin by sampling the SMPL-X parameters from large-scale human motion capture datasets such as AMASS [27]. Then, we utilize a random camera setup to render the human mesh into three spatial maps, i.e., a keypoint heatmap, a depth map, and a surface normal map. These spatial maps are added as extra input conditions to the generative models, as they provide valuable information about human body’s pose, depth, and surface orientation, enhancing the model’s ability to generate more accurate outputs. Finally, we feed text prompts and these perceptual maps to a multi-condition ControlNet [28] for generating human images.
Notably, we observe existing general-purpose controllable image generators, e.g., ControlNet [28], struggle to generate accurate human images even though precise spatial conditioning controls are provided. This is mainly due to the lack of high-quality training samples. To address this, we first create a large-scale human-centric dataset, each with detailed annotations including text descriptions, human poses, depth maps, and surface normals. The full data curation pipeline is outlined in Sec.˜3.2. We then fine-tune a multi-condition ControlNet using this dataset to improve its performance in generating human images under spatial constraints.
Thanks to the high-quality fine-tuning data, most data samples generated by the multi-condition ControlNet demonstrate satisfactory performance. However, empirical analysis revealed instances of label noise in the initial training pairs. These issues include cases where the generated human figure and input conditions appear as mirrored pairs (see  Fig.˜6), or where the head orientation in the image differs from the input SMPL-X parameters (see failure cases in Supp. Mat.). To address this, we use a pre-trained segmentation model, SAM [29], to predict the human mask in the generated images. We then calculate the intersection-over-union (mIoU) between the ground-truth human mask and the predicted one, to filter out data samples with an IoU below a certain threshold. Finally, we apply a well-trained 2D dense surface estimator to refine the SMPL-X parameters using the SMPLify method [16], further improving the alignment between the input and generated outputs.
@@
With the aforementioned pipeline, we can finally generate a large-scale 3D human dataset in the wild, with high-quality human appearance and 3D annotations. As shown in Table˜I, both mocap-based and CG-rendered datasets are limited in providing diverse human identities and in-the-wild scenes. Compared to previous datasets, our pipeline can generate diverse human identities, human interactions, and various in-the-wild scenes. Notably, the pipeline is much cheaper than both mocap-based and CG-based counterparts and is scalable to generate 3D human datasets in the wild with versatile real-world scenes.
Data type Datasets Subjects Scene Frame Modalities
RGB D/N K2D K3D B.P. WB.P.
Real-world Monocular COCO [21] - - 104K
MPI-INF-3DHP [15] 8 - 1.4M
MPII [22] - - 24K
Multi-view HuMMan 1000 1 60M
ZJU Mocap [30] 6 1 >1K
AIST++ [31] 30 1 10.1M
MOCAP 3DPW [23] 5 <60 51K
Human3.6M [12] 11 1 3.6M
Synthetic AGORA [32] >350 - 18K
Synbody [20] 10K 6 1.2M
BEDLAM [19] 217 103 280K
Generated HumanWild $\infty$ $\infty$ 630K
TABLE I: Comparison of different types of 3D HPS datasets.
@@
Our contributions can be summarised as follows. 1). We curate a large-scale, human-centric dataset with comprehensive, high-quality annotations designed for controllable human image generation. 2). Without pursuing fancy techniques, we propose a simple yet cost-efficient pipeline to synthesize realistic and diverse human images with well-aligned annotations, including paired SMPL-X parameters and human images. Beyond human pose and shape estimation, the dataset can empower a wide range of downstream perception tasks by rendering SMPL-X mesh into corresponding annotation format, e.g., human part segmentation. 3). We verify the quality of the generated dataset on the 3D HPS task. Experimental results indicate that the proposed pipeline is compositional with CG-rendered data, enhancing performance across multiple challenging HPS benchmarks under consistent settings.
2 Related Work
@@ 2.1 Human Pose and Shape Estimation Datasets
Synthetic human pose datasets, developed with computer graphics techniques, have been used for many years. SURREAL [36] applies human skin and cloth textures to bare SMPL meshes, which lack realistic details. AGORA [32] uses high-quality static human scans for image rendering, but this routine also suffers from a high workload of scanning and rigging. However, rendering realistic, manipulable synthetic human datasets involves many challenges, including the need for diverse virtual properties for realistic data. BEDLAM [19] and Synbody [20] augment SMPL-X meshes [37] with diverse hair models and skin textures, facilitating the simulation of physically realistic clothes and hair dynamics. These processes can be resource-intensive. Furthermore, the use of rendering engines demands many professional skills. Besides, the rendering process can be computationally expensive and time-consuming.
Controllable human image generation has gained great traction with the advancement of Stable Diffusion [38, 28]. Text2Human [39] uses a diffusion-based transformer sampler in response to text prompts and predicts indices from a hierarchical texture-aware codebook to conditionally generate realistic human images. HumanSD [40] introduces a skeleton-guided diffusion model with a novel heatmap loss for pose-conditioned human image generation. HyperHuman [41] proposes to jointly denoise surface normal and depth along with the synthesized RGB image conditioned with text prompt and pose skeleton.
Generative Models for Perception Tasks is a rapidly growing field that aims to enhance the performance of perception models by utilizing the vast prior knowledge of generative models. Several pioneering works have proposed generating perception datasets with diffusion models. For instance, [42] and [43] verified the effectiveness of datasets synthesized by generative models in fundamental perception tasks, i.e., image classification, and object detection. StableRep [44] argues that training modern self-supervised methods on synthetic images from Stable Diffusion Models can yield impressive results. The learned representations often surpass those learned from real images of the same sample size. DatasetDM [45] leverages the prior of the Stable Diffusion model by training customized perception decoders upon the output of the UNet [46] and then employs it to generate pseudo labels for semantic segmentation, depth estimation, and 2D human pose estimation.
@@ 3 Method
Figure 2: The overall pipeline of the proposed controllable data generation. Our ControlNet could be conditioned on fine-grained keypoint maps, depth maps, surface normal maps, and structural text prompts. Our text prompt includes human appearance, pose, and indoor/outdoor scene types based on w. or w/o background normals.
We present HumanWild, a simple yet effective approach for creating versatile human body images and corresponding 3D parametric mesh annotations in a fully automated fashion, which can be used for many downstream human perception tasks, such as 2D/3D human pose and shape estimation (see Fig.˜2). The core idea of the proposed pipeline is creating large-scale image-mesh-caption pairs by incorporating 2D generative models, e.g., ControlNet [28] and 3D human parametric models [37]. For the sake of completeness, we give a brief review of the controllable text-to-image (T2I), image-to-image (I2I) generative models, and the 3D human parametric model, SMPL-X [37] in Sec.˜3.1. Then, we show how we build large-scale human-centric datasets in Sec.˜3.2 for the controllable human generation task. Next, we illustrate how we generate the initial human image-annotation pairs in Sec.˜3.3. Furthermore, we show our filtering strategy for the noisy labels to get high-quality training pairs in Sec.˜3.4.
@@ 3.1 Prerequisites
Stable Diffusion Models [38, 24] are text-to-image diffusion models capable of generating highly photo-realistic images given any text input. It consists of an autoencoder and a denoiser. During training, U-Net model is designed to take a textual description and noise map $\epsilon$ as input, learning a forward noising process by gradually transforming an image latent variable, denoted as $\bar{y}\in\mathbb{R}^{h\times w\times c}$ into a noise map $\bar{y_{t}}\in\mathbb{R}^{h\times w\times c}$:
@@
where $t=[1,\cdots,T]$ is the timestep for controlling the noise level and ${\alpha}_{t}$ is the noise scheduler [47], $h$, $w$ and $c$ are the height, width, and channels of the latent variable. During inference, UNet takes a noise map as input and generates an approximation of the original latent variable by using a denoise scheduler [48].
@@
SMPL-X [37], defined as $M(\boldsymbol{\beta},\boldsymbol{\theta},\boldsymbol{\psi}):\mathbb{R}^{|\theta|\times|\beta|\times|\psi|}\rightarrow\mathbb{R}^{3N}$, is a 3D wholebody human parametric model, employing shape $\boldsymbol{\beta}\in\mathbb{R}^{200}$, expression $\boldsymbol{\psi}\in\mathbb{R}^{50}$, and pose $\boldsymbol{\theta}\in\mathbb{R}^{55\times 3}$ to control the entire body mesh. The function of SMPL-X provides a differentiable skinning process that uses pose, shape, and expression parameters as inputs and delivers a triangulated mesh $V\in\mathbb{R}^{N\times 3}$ with $N=10475$ vertices. The reconstructed 3D joints $J\in\mathbb{R}^{144\times 3}$ can be obtained using a forward kinematics process. We use the SMPL-X model to represent the human body.
3.2 Human-centric Dataset for Human Generation
Large-scale datasets with high-quality samples, rich annotations, and diverse distributions are essential for controllable human image generation. To address this need, we have created a comprehensive human dataset with extensive annotations. Our dataset consists of two parts: CG-rendered human images and curated human images from existing sources, such as LAION-Human [49] and other publicly available Internet datasets.
Real-world Human. For real-world human data collection, we filter images from LAION-2B-en [49] and COYO-700 [50] using the YOLOS human detector [51]. Next, we apply state-of-the-art perception models—DepthAnything-V2 [52], DSINE [53], and DWPose [54]—to generate corresponding depth maps, surface normal maps, and keypoint heatmaps, respectively. The curated dataset covers a broad spectrum of scenes, including diverse backgrounds and partial human regions like clothing and limbs, providing a rich variety of real-world data for further use.
CG-rendered Human. We also introduce a synthetic dataset to provide accurate annotations for controllable human image generation. Specifically, we utilize 630 human models from RenderPeople [55] and 660 body pose sequences from AMASS [27]. We use HDRi images with various lighting conditions as backgrounds. As shown in Fig.˜3, the rendered dataset includes diverse human identities along with precise depth maps and surface normal maps without background, enhancing the dataset’s utility for various tasks.
Figure 3: Visualization of our rendered human dataset with diverse human identities and in-the-wild scenes. Each data sample contains RGB image (left), a depth map (middle), and a surface normal map (right).
Figure 4: Visualization of HPS estimation. For each data sample, left is the original image, middle is trained on BEDLAM dataset, right is finetuned on our HumanWild dataset.
Fine-tune Multi-condition ControlNet. We then use the curated dataset for fine-tuning a multi-condition ControlNet [28] based on SDXL [24] for our customized controllable human image generation. The input conditions of ControlNet are text prompts, keypoint heatmaps, depth maps, and surface normal maps. As shown in Fig.˜5, our fine-tuned ControlNet are generalizable to diverse poses and shapes generated by SMPL-X, while existing ControlNet variants trained on general datasets failed in many scenes.
Figure 5: Visualization of human images generated by our multi-condition generation pipeline with diverse poses and scenes. For each data sample, the first column is generated from the ControlNet before revision. The second and third columns are generated from our revised model. The fourth, fifth, and sixth columns are keypoint heatmap, depth map, and surface normal map rendered from the SMPL-X model.